# Audio–Visual Speech Enhancement: Architectural Design and Deployment Strategies

### v2

@@ IV-A Algorithm Processing Latency
than the duration of a frame, thus reducing the user experience latency to be within $100$ ms. However, reducing the input chunk size will result in a serious performance loss. Figure 6 plots the time-domain waveforms and spectrogram of both the noisy input audio and the output audio post-enhancement. The visualization is stratified by varying lengths of input chunks, clearly showcasing the interplay between chunk size and signal integrity. Speech enhanced with $10$s audio and video inputs yields a distinct and unblemished vocal output from the target speaker. When the enhancement is applied on a frame-by-frame basis, the resulting output is predominantly tainted with noise. A notable improvement in audio clarity is observed as we extend the duration of the input; however, this does not uniformly guarantee optimal performance across all temporal segments. Some intervals exhibit satisfactory enhancement, while others remain suboptimal. Consequently, to achieve a consistently high Quality of Service, we found it necessary to utilize 10-second blocks of audio and video for each processing iteration, which resulted in about $2.2$s of algorithm processing time, and to make sure that the server could complete each processing, we set an interval of 2.32.42.3-2.4s, which resulted in a total latency of about $5$s.
@@
IV-B Algorithm development
To mitigate the inherent latency in our system, we tried to refine our network’s proficiency in distilling potent feature representations from short segments of audio and visual data. To this end, we have calibrated the network training scheme to work with condensed input durations of merely $200$ ms. Leveraging the architecture depicted in Figure 7, we constructed a series of networks, reducing their structures by paring down layers and neural nodes with the express aim of shortening processing delays. Table I lists the total parameters and processing latency of each of the three networks.
Figure 7: Audio-Visual Speech Enhancement AI model architecture.
Models Total parameters Time complexity
Model 1 1,540,396 (5.88 MB) $\tilde{1}200$ ms
Model 2 603,564 (2.30 MB) $\tilde{5}50$ ms
Model 3 202,564 (791.27 KB) $\tilde{3}50$ ms
TABLE I: MODELS’ PARAMETERS AND COMPLEXITY
Figure 8 plots the time-domain waveforms and spectrogram of the original target speaker’s voice, noisy input audio, and the output audio post-enhancement. The reduction in visual information due to the shortened input duration results in a lower reproduction performance of the original sound when it is weak, but the noise reduction is still satisfactory.
@@ IV-B Algorithm development
V CONCLUSION
This study demonstrated a complete 5G cloud-based audio-visual speech enhancement (AVSE) system that integrates CNN-based acoustic processing and OpenCV facial feature extraction with an LSTM fusion network for temporal coherence. Experimental results showed that real-time performance depends critically on network latency and algorithmic processing time. Only 5G and wired Ethernet consistently met the coherence requirement ($t_{\text{comm}}\leq t_{\text{chunk}}$) (40 ms) for uncompressed 0.3 MB audio-video chunks, while compression at a factor of 80 reduced payloads to about 70 KB with negligible degradation, enabling smoother operation in constrained networks.
The findings highlight that network quality, chunk size, and model complexity jointly determine system responsiveness and enhancement quality. Although shorter input windows and smaller models reduce processing latency—from about 1.2 s to 0.35 s—they also degrade speech reconstruction, particularly under weak target signals. Achieving optimal performance therefore requires balancing temporal context and computational efficiency. Future work will focus on adaptive chunking, lightweight streaming architectures, and edge-assisted processing to further minimize delay while preserving enhancement fidelity in real-world wireless environments.
@@ Acknowledgements
This work was supported by the UK Engineering and Physical Sciences Research Council (EPSRC) Grant No. EP/XXXXXXX/X. The authors would like to gratefully acknowledge Haifeng Luo from the University of XXXX for his contribution in implementing XXXX.

### v4

@@ V-B Algorithm Processing Latency
The end-to-end user-perceived delay ($t_{\text{delay}}$) can be decomposed into communication latency ($t_{\text{comm}}$) and algorithm processing latency ($t_{\text{proc}}$), i.e.,
$t_{\text{delay}}=t_{\text{comm}}+t_{\text{proc}}.$ (12)
In practice, $t_{\text{comm}}$ mainly affects the synchronization and coherence of the feedback audio, whereas $t_{\text{proc}}$ directly determines the system responsiveness. Real-world measurements indicate that modern wireless standards such as 5G NR and Wi-Fi 6 can satisfy the required throughput and transmission delay constraints. Consequently, minimizing $t_{\text{proc}}$ becomes the dominant factor in reducing overall interaction latency.
A common approach to reduce $t_{\text{proc}}$ is to shorten the input chunk duration and reduce model complexity. However, in audio–visual speech enhancement (AVSE), the enhancement model must capture temporal lip-motion dynamics in order to reliably extract the target speaker from competing noise sources. As a result, overly short input chunks may degrade enhancement performance due to insufficient temporal context.
This latency–performance trade-off has been widely recognized in the AVSE literature. Gogate et al. introduced CochleaNet, a causal, language-, noise-, and speaker-independent AVSE framework designed for real-time processing in noisy environments [5]. More recently, Gogate et al. proposed a robust real-time AVSE framework that integrates a CycleGAN-based visual enhancement stage with a causal DNN-based fusion model to estimate an ideal binary mask (IBM) for noise suppression [7]. Their system reports an end-to-end processing latency of approximately $15$ ms while achieving significant gains in objective and subjective intelligibility metrics, demonstrating the feasibility of low-latency AVSE under real-time constraints [7]. Additional efforts by the same group further explore lightweight architectures for benchmark AVSE challenges, emphasizing computational efficiency and practical deployment [6]. Furthermore, latency-efficient AVSE models based on cross-modal feature synchronization have reported real-time performance with processing delays on the order of $36$ ms, reinforcing the importance of low-latency multimodal fusion for assistive hearing applications [11].
In our system, the input stream is segmented into chunks of $N$ frames, where each frame corresponds to $40$ ms. Fig. 8 illustrates the measured algorithm processing latency as a function of chunk size. The results indicate that processing latency increases approximately linearly with the input duration. In principle, processing each frame independently could reduce $t_{\text{proc}}$ below $40$ ms, enabling a total interaction latency within $100$ ms. However, such short-context processing significantly reduces enhancement quality due to the lack of sufficient temporal information for robust audio–visual fusion.
To further evaluate the impact of chunk size on enhancement quality, Fig. 9 shows the time-domain waveform and spectrogram of the noisy input and enhanced output audio under varying chunk durations. Using $10$-s audio–visual segments produces the strongest reconstruction of the target speech and the most effective suppression of competing background noise. In contrast, frame-level enhancement yields output dominated by residual noise, indicating that the model fails to extract reliable speaker-discriminative features at this temporal resolution. Increasing the chunk duration generally improves enhancement quality, although the performance is not uniformly consistent across all temporal segments, suggesting that longer temporal context does not necessarily guarantee optimal reconstruction in all intervals.
@@
V-C Algorithm Development
To mitigate the processing latency while maintaining enhancement performance, we investigated whether the proposed model could learn discriminative representations from shorter temporal contexts. Specifically, the training scheme was adapted to support condensed input durations of $200$ ms. Based on the AVSE architecture shown in Figure 1, a series of lightweight variants were developed by reducing network depth and neuron count with the objective of shortening inference time. The total number of parameters and measured execution time for each reduced model are summarized in Table III.
TABLE III: Model Complexity and Execution Time
Model Total Parameters Execution Time (ms)
Model 1 1,540,396 (5.88 MB) 1200
Model 2 603,564 (2.30 MB) 550
Model 3 202,564 (791.27 KB) 350
Fig. 10 compares the time-domain waveform and spectrogram of the clean target speech, noisy mixture, and enhanced output produced by the reduced models. The results show that shortening the input duration and reducing model capacity decreases reconstruction fidelity, particularly for low-energy speech components, due to reduced availability of visual temporal information. Nevertheless, the reduced models maintain satisfactory noise suppression performance, indicating that lightweight architectures can provide an effective trade-off between enhancement quality and low-latency execution.
@@ V-C Algorithm Development
Overall, the results demonstrate a clear trade-off between processing latency and enhancement performance. Longer input chunks provide richer temporal context for lip-motion tracking and speaker discrimination, but significantly increase processing delay. Conversely, reducing chunk size and model complexity improves responsiveness but may degrade speech reconstruction quality in challenging low-SNR conditions. Therefore, selecting an optimal configuration requires balancing the latency constraints of real-time feedback against the temporal context required for robust AVSE.
VI CONCLUSION
This work presented a comprehensive evaluation of a real-time, cloud–edge–assisted audio-visual speech enhancement (AVSE) system operating over a public 5G edge cloud. The proposed system integrates CNN-based acoustic processing and OpenCV-driven facial feature extraction with an LSTM fusion network to preserve temporal coherence, and was deployed and evaluated using a Vodafone-compatible AWS Wavelength edge in London. The results demonstrate that compute placement at the network edge and network transport latency are primary determinants of end-to-end performance for interactive multimedia and perceptual enhancement applications.
Experimental analysis showed that real-time coherence is achieved only when communication delay remains below the chunk duration ($t_{\text{comm}}\leq t_{\text{chunk}}=40$ ms). In practice, this requirement was consistently met by 5G and wired Ethernet connections for uncompressed 0.3 MB audio-video chunks, while other access networks failed to maintain stable operation. Applying aggressive compression (up to a factor of 80) reduced payloads to approximately 70 KB with negligible perceptual degradation, significantly improving robustness under constrained uplink conditions and enabling smoother real-time streaming. These findings confirm that uplink capacity is often the dominant bottleneck for interactive edge-based multimedia services.
We further observed that system responsiveness and enhancement quality are jointly governed by network quality, chunk size, and model complexity. Reducing temporal context and model size lowered processing latency from approximately 1.2 s to 0.35 s, but at the cost of degraded speech reconstruction, particularly in low-SNR conditions. This reveals a fundamental trade-off between computational efficiency and enhancement fidelity, underscoring the need for adaptive operating profiles that can dynamically balance latency, throughput, and quality based on current network and compute conditions. Stress testing identified safe operating regions as well as the importance of fallback modes, such as audio-only operation, to preserve service continuity during adverse network conditions.
Overall, the results indicate that public 5G edge environments can sustain real-time, interactive multimedia workloads, provided that network and compute resources are carefully orchestrated. However, performance margins remain tighter than in dedicated or private infrastructures, making dynamic adaptation and edge proximity essential architectural considerations. The insights derived from this study are directly applicable to delay-sensitive services such as augmented reality, interactive conferencing, and multimodal perception enhancement.
Future work will extend this study by incorporating mobility and multi-user scenarios, exploring GPU-accelerated edge instances for more complex models, and evaluating performance across multiple geographic regions and operator ecosystems. We also plan to investigate adaptive chunking, lightweight streaming architectures, and edge-assisted learning to further reduce latency while preserving enhancement fidelity in real-world wireless environments.
@@ Acknowledgements
This work was supported by the UK Engineering and Physical Sciences Research Council (EPSRC) Grant No. EP/T021063/1.